### v1

# On Strongly Regular Graphs and the Friendship Theorem

Abstract. This paper presents an alternative proof of the celebrated friendship theorem, originally established by Erdős, Rényi, and Sós (1966). The proof is based on a closed-form expression for the Lovász $\vartheta$-function of strongly regular graphs, which was recently derived by the author. Leveraging this expression, the paper also establishes new necessary conditions for a strongly regular graph to be a spanning subgraph or an induced subgraph of another strongly regular graph. In the case of induced subgraphs, the analysis further incorporates a property of graph energies.
Keywords. Friendship theorem, strongly regular graph, Lovász $\vartheta$-function, graph invariants, graph energy, windmill graph, spanning subgraph, induced subgraph.
2020 Mathematics Subject Classification (MSC). 05C50, 05C60, 05E30.
1 Introduction
The friendship theorem in graph theory states that if a finite graph has the property that every pair of distinct vertices shares exactly one common neighbor, then the graph must be a collection of triangles that share a common vertex. In other words, the theorem asserts that such a finite graph is necessarily a friendship graph (also known as a windmill graph), meaning that it consists of a central vertex connected to all others with additional edges forming triangles. The name of this theorem stems from its human interpretation, where if every two individuals in a group of people have exactly one mutual friend, then there must be someone in that group who is everybody’s friend (often referred to as the ”politician”).
That theorem was first proved by Erdős, Rényi, and Sós [5], and it found diverse applications in combinatorics and other disciplines. While the friendship theorem itself is a result in graph theory, its structural properties make it useful in designing graph-based codes, network coding, and cryptographic schemes. In network coding, graphs represent information flow between nodes, and the friendship property ensures that all transmissions go through a single relay node, which models centralized communication structures such as satellite communication, where every ground station interacts via a single hub. This property is also useful in wireless sensor networks, where multiple sensors communicate through a central processing node. Furthermore, friendship graphs can be linked to block designs and combinatorial structures used in constructing error-correcting codes with good covering properties. Beyond coding, the friendship theorem has a broad range of applications in networks, biology, computer science, social sciences, and economics. The fact that every two vertices share exactly one mutual neighbor makes it a useful model for systems where a single hub controls multiple independent interactions, such as in social influence, network protocols, and centralized coordination problems.
The original proof of the theorem involves arguments based on graph connectivity, degree counting, and combinatorial reasoning. The reader is referred to a nice exposition of a proof of the friendship theorem in Chapter 44 of [1], which combines combinatorics and linear algebra (spectral graph theory). Due to its importance and simplicity, the theorem has been proved in various ways. The reader is also referred to [13] for historical notes on its proofs, as well as a new combinatorial proof proposed therein.
The aim of this paper is to provide a different proof of that fundamental theorem, which relies on a recent finding of the author [14], where a closed-form expression of the celebrated Lovász $\vartheta$-function has been derived for the structured family of strongly regular graphs. We believe that this proof offers a new perspective to the available collection of proofs of the friendship theorem. Leveraging this expression, the paper also establishes new necessary conditions for a strongly regular graph to be a spanning subgraph or an induced subgraph of another strongly regular graph. In the case of induced subgraphs, the analysis further incorporates a property of graph energies. In general, understanding whether a graph is a spanning subgraph or an induced subgraph of another given graph is a meaningful and widely applicable question across theoretical, algorithmic, and applied fields of graph theory.
This paper is structured as follows. Section 2 provides some essential preliminaries for our alternative proof of the friendship theorem in Section 3. A discussion on a spectral graph-theoretic proof that is closely related to the proof presented in [1, Chapter 44] is also provided in Section 3. Based on the Lovász $\vartheta$-function of strongly regular graphs, Sections 4 and 5 establish new necessary conditions for one strongly regular graph to be either a spanning subgraph or an induced subgraph, respectively, of another strongly regular graph. The proposed conditions can be easily checked, and their utility is exemplified.
2 Preliminaries
Definition 2.1 (Lovász $\vartheta$-function, [11]).
Let $\mathsf{G}$ be a finite, undirected and simple graph. The Lovász $\vartheta$-function of $\mathsf{G}$ is defined as
$\vartheta(\mathsf{G})\triangleq\min_{\{\bf{u}_{i}\},\bf{c}}\,\max_{i\in\operatorname{\mathsf{V}}(\mathsf{G})}\,\frac{1}{\bigl{(}{\bf{c}}^{\mathrm{T}}{\bf{u}}_{i}\bigr{)}^{2}}\,,$ (1)
where the minimum is taken over all orthonormal representations $\{{\bf{u}}_{i}:i\in\operatorname{\mathsf{V}}(\mathsf{G})\}$ of $\mathsf{G}$, and over all unit vectors ${\bf{c}}$. The unit vector $\bf{c}$ is called the handle of the orthonormal representation. By the Cauchy-Schwarz inequality,
$\bigl{|}{\bf{c}}^{\mathrm{T}}{\bf{u}}_{i}\bigr{|}\leq\|{\bf{c}}\|\,\|{\bf{u}}_{i}\|=1\;\Longrightarrow\;\vartheta(\mathsf{G})\geq 1,$
with equality if and only if $\mathsf{G}$ is a complete graph.
Without loss of generality, the dimension $d$ can be set to be equal to the order of $\mathsf{G}$, i.e., $d=|\operatorname{\mathsf{V}}(\mathsf{G})|$.
Let
  • ${\bf{A}}$ is the $n\times n$ adjacency matrix of $\mathsf{G}$   ($n\triangleq|\operatorname{\mathsf{V}}(\mathsf{G})|$);
  • ${\bf{J}}_{n}$ is the all-ones $n\times n$ matrix;
  • ${\bf{\mathcal{S}}}_{+}^{n}$ is the set of all $n\times n$ positive semidefinite matrices.
The following semidefinite program (SDP) can be used for computing $\vartheta(\mathsf{G})$ (see [11] and [12, Chapter 11]):
$\begin{array}[]{l}\text{maximize}\;\;\text{Trace}({\bf{B}}\,{\bf{J}}_{n})\\ \text{subject to}\\ \begin{cases}{\bf{B}}\in{\bf{\mathcal{S}}}_{+}^{n},\;\;\text{Trace}({\bf{B}})=1,\\ A_{i,j}=1\;\Rightarrow\;B_{i,j}=0,\quad i,j\in[n].\end{cases}\end{array}$ (2)
This renders the computational complexity of $\vartheta(\mathsf{G})$ feasible. More explicitly, there exist some standard algorithms in convex optimization that numerically compute $\vartheta(\mathsf{G})$, for every graph $\mathsf{G}$, with precision of $r$ decimal digits, and in polynomial-time in $n$ and $r$.
Let $\alpha(\mathsf{G})$, $\operatorname{\omega}(\mathsf{G})$, and $\operatorname{\chi}(\mathsf{G})$ denote the independence number, clique number, and chromatic number of a graph $\mathsf{G}$. Then, the Lovász $\vartheta$-function of $\mathsf{G}$ and its complement graph $\overline{\mathsf{G}}$ satisfy the following properties (see [11, 12]):
  1. (1)
    Sandwich theorem:
    $\alpha(\mathsf{G})\leq\vartheta(\mathsf{G})\leq\operatorname{\chi}(\overline{\mathsf{G}}),$ (3)
    $\operatorname{\omega}(\mathsf{G})\leq\vartheta(\overline{\mathsf{G}})\leq\operatorname{\chi}(\mathsf{G}).$ (4)
  2. (2)
    Computational complexity: The graph invariants $\alpha(\mathsf{G})$, $\operatorname{\omega}(\mathsf{G})$, and $\operatorname{\chi}(\mathsf{G})$ are NP-hard problems. However, as mentioned above, the numerical computation of $\vartheta(\mathsf{G})$ is in general feasible.
  3. (3)
    Hoffman-Lovász inequality: Let $\mathsf{G}$ be $d$-regular of order $n$. Then,
    $\vartheta(\mathsf{G})\leq-\frac{n\,\lambda_{n}(\mathsf{G})}{d-\lambda_{n}(\mathsf{G})},$ (5)
    with equality if $\mathsf{G}$ is edge-transitive.
Definition 2.2 (Strongly Regular Graphs).
Let $\mathsf{G}$ be a $d$-regular graph of order $n$. The graph $\mathsf{G}$ is a strongly regular graph (SRG) if there also exist nonnegative integers $\lambda$ and $\mu$ such that
  • Every pair of adjacent vertices have exactly $\lambda$ common neighbors;
  • Every pair of distinct and non-adjacent vertices have exactly $\mu$ common neighbors.
Such a strongly regular graph is said to be a graph within the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$.
It should be noted that within the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$, with fixed four parameters $n$, $d$, $\lambda$, and $\mu$, there may exist several strongly regular graphs that are nonisomorphic to each other. For example, according to [3], there exist 167 nonisomorphic strongly regular graphs within the family $\operatorname{\mathsf{srg}}(64,18,2,6)$.
Theorem 2.1 (Bounds on Lovász function of Regular Graphs, [14]).
Let $\mathsf{G}$ be a $d$-regular graph of order $n$, which is a non-complete and non-empty graph. Then, the following bounds hold for the Lovász $\vartheta$-function of $\mathsf{G}$ and its complement $\overline{\mathsf{G}}$:
  1. (1)
    $\frac{n-d+\lambda_{2}(\mathsf{G})}{1+\lambda_{2}(\mathsf{G})}\leq\vartheta(\mathsf{G})\leq-\frac{n\lambda_{n}(\mathsf{G})}{d-\lambda_{n}(\mathsf{G})}.$ (6)
    • Equality holds in the leftmost inequality of (6) if $\overline{\mathsf{G}}$ is both vertex-transitive and edge-transitive, or if $\mathsf{G}$ is a strongly regular graph;
    • Equality holds in the rightmost inequality of (6) if $\mathsf{G}$ is edge-transitive, or if $\mathsf{G}$ is a strongly regular graph.
  2. (2)
    $1-\frac{d}{\lambda_{n}(\mathsf{G})}\leq\vartheta(\overline{\mathsf{G}})\leq\frac{n\bigl{(}1+\lambda_{2}(\mathsf{G})\bigr{)}}{n-d+\lambda_{2}(\mathsf{G})}.$ (7)
    • Equality holds in the leftmost inequality of (7) if $\mathsf{G}$ is both vertex-transitive and edge-transitive, or if $\mathsf{G}$ is a strongly regular graph;
    • Equality holds in the rightmost inequality of (7) if $\overline{\mathsf{G}}$ is edge-transitive, or if $\mathsf{G}$ is a strongly regular graph.
As a common sufficient condition, note that all the inequalities in (6) and (7) hold with equality if $\mathsf{G}$ is a strongly regular graph. (Recall that a graph $\mathsf{G}$ is strongly regular if and only if its complement $\overline{\mathsf{G}}$ is so). The following result provides a closed-form expression of the Lovász $\theta$-function of all strongly regular graphs.
Corollary 2.1 (The Lovász $\vartheta$-Function of Strongly Regular Graphs, [14]).
Let $\mathsf{G}$ be a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$. Then,
$\vartheta(\mathsf{G})=\dfrac{n\,(t+\mu-\lambda)}{2d+t+\mu-\lambda},$ (8)
$\vartheta(\overline{\mathsf{G}})=1+\dfrac{2d}{t+\mu-\lambda},$ (9)
where
$t\triangleq\sqrt{(\mu-\lambda)^{2}+4(d-\mu)}.$ (10)
Remark 2.1.
In light of Corollary 2.1, all strongly regular graphs within a family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ are not only cospectral graphs [2] but they also share the same value of the Lovász $\vartheta$-function. In that context, it should be noted that there exist connected, cospectral, and irregular nonisomorphic graphs with an identical vertex set that are jointly cospectral with respect to several of their associated matrices (including the adjacency matrix) and share identical values of certain graph invariants, such as the independence number, clique number, and chromatic number, and yet they differ in their Lovász $\vartheta$-functions. The reader is referred to [15] for such explicit graph constructions and related analysis.
Corollary 2.2 (A Relation for Strongly Regular Graphs, [14]).
The relation
$\vartheta(\mathsf{G})\,\vartheta(\overline{\mathsf{G}})=n,$ (11)
holds not only for all vertex-transitive graphs (see [11, Theorem 8]), but it also holds for all strongly regular graphs (in general, strongly regular graphs may not be vertex-transitive, see, e.g., [6]).
3 The Friendship Theorem: An Alternative Proof
This paper presents an alternative proof of the celebrated friendship theorem, originally established in [5], which relies on the Lovász $\vartheta$-function of strongly regular graphs in Corollary 2.1.
Theorem 3.1 (Friendship Theorem, [5]).
Let $\mathsf{G}$ be a finite graph in which any two distinct vertices have a single common neighbor. Then, $\mathsf{G}$ has a vertex that is adjacent to every other vertex.
A Human Interpretation of Theorem 3.1 is well known. Assume that there is a party with $n$ people, where every pair of individuals has precisely one common friend at the party. Theorem 3.1 asserts that one of these people is a friend of everyone. Indeed, construct a graph whose vertices represent the $n$ people, and every two vertices are adjacent if and only if they represent two friends. The claim then follows from Theorem 3.1.
$u$
Figure 1: Windmill graph (also known as friendship graph) consisting of multiple triangles sharing a common central vertex $u$.
Remark 3.1 (On Theorem 3.1).
The windmill graph (see Figure 1) has the desired property of the friendship theorem and turns out to be the only graph with that property. Remarkably, the friendship theorem does not hold for infinite graphs. Indeed, for an inductive construction of a counterexample, one may start with a 5-cycle $\operatorname{\mathsf{C}}_{5}$, and repeatedly add a common neighbor for every pair of vertices that does not yet have one. This process results in a countably infinite friendship graph without a vertex adjacent to all other vertices.
Next, we present an alternative proof of Theorem 3.1, based on our expression in Corollary 2.1 for the Lovász $\vartheta$-function of strongly regular graphs and their complements, which are also strongly regular.
Proof.
Suppose the assertion is false, and let $\mathsf{G}$ be a counterexample — a finite graph in which any two distinct vertices have a single common neighbor, yet no vertex in $\mathsf{G}$ is adjacent to all other vertices. A contradiction is obtained as follows:
The first step shows that the graph $\mathsf{G}$ is regular, as proved in Chapter 44 of [1]. We provide a variation of that regularity proof, and then take a different approach, relying on graph invariants such as $\operatorname{\omega}(\mathsf{G}),\operatorname{\chi}(\mathsf{G})$, and $\vartheta(\mathsf{G})$.
  • To assert the regularity of $\mathsf{G}$, it is first proved that nonadjacent vertices in $\mathsf{G}$ have equal degrees, i.e., $d(u)=d(v)$ if $\{u,v\}\not\in\operatorname{\mathsf{E}}(\mathsf{G})$.
  • The given hypothesis yields that $\mathsf{G}$ is a connected graph. Let $\{u,v\}\not\in\operatorname{\mathsf{E}}(\mathsf{G})$, and let $\mathcal{N}(u)$ and $\mathcal{N}(v)$ denote, respectively, the sets of neighbors of the nonadjacent vertices $u$ and $v$.
  • Let $f\colon\mathcal{N}(u)\to\mathcal{N}(v)$ be the injective function where every $x\in\mathcal{N}(u)$ is mapped to the unique $y\in\mathcal{N}(x)\cap\mathcal{N}(v)$. Indeed, if $z\in\mathcal{N}(u)\setminus\{x\}$ satisfies $f(z)=y$, then $x$ and $z$ share two common neighbors (namely, $y$ and $u$), which contradicts the assumption of the theorem.
  • Since $f\colon\mathcal{N}(u)\to\mathcal{N}(v)$ is injective, it follows that $|\mathcal{N}(u)|\leq|\mathcal{N}(v)|$.
  • By symmetry, swapping $u$ and $v$ (as nonadjacent vertices) also yields $|\mathcal{N}(v)|\leq|\mathcal{N}(v)|$, so we get $d(u)=|\mathcal{N}(u)|=|\mathcal{N}(v)|=d(v)$ for all vertices $u,v\in\operatorname{\mathsf{V}}(\mathsf{G})$ such that $\{u,v\}\not\in\operatorname{\mathsf{E}}(\mathsf{G})$.
  • To complete the proof that $\mathsf{G}$ is regular, let $u$ and $v$ be nonadjacent vertices in $\mathsf{G}$. By assumption, except of one vertex, all vertices are either nonadjacent to $u$ or $v$. Hence, except of that vertex, all these vertices must have identical degrees by what we already proved.
  • Finally, by our further assumption (later leading to a contradiction), since there is no vertex in $\mathsf{G}$ that is adjacent to all other vertices, also the single vertex that is adjacent to $u$ and $v$ has a nonneighbor in $\mathsf{G}$, so it also should have an identical degree to all the degrees of the other vertices by what is proved above. Consequently, $\mathsf{G}$ is a regular graph.
From that point, our proof proceeds differently.
  • Let $\mathsf{G}$ be a $k$-regular graph on $n$ vertices. By assumption, every two vertices have exactly one common neighbor, so $\mathsf{G}$ is a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,k,1,1)$.
  • First, if $k=1$ or $k=2$, then by assumption, it follows that $\mathsf{G}=\operatorname{\mathsf{K}}_{1}$ or $\mathsf{G}=\operatorname{\mathsf{K}}_{2}$, respectively, leading to a contradiction. Hence, let $k\geq 3$.
  • Two adjacent vertices in $\mathsf{G}$ share a common neighbor, so $\mathsf{G}$ contains a triangle. Moreover, $\mathsf{G}$ is $\operatorname{\mathsf{C}}_{4}$-free since every two vertices have exactly one common neighbor, so it is $\operatorname{\mathsf{K}}_{4}$-free. Hence, $\operatorname{\omega}(\mathsf{G})=3$.
  • We next show that $\operatorname{\chi}(\mathsf{G})=3$. First, $\operatorname{\chi}(\mathsf{G})\geq\operatorname{\omega}(\mathsf{G})=3$. We also need to show that $\operatorname{\chi}(\mathsf{G})\leq 3$, which means that three colors suffice to color the vertices of $\mathsf{G}$ in a way that no two adjacent vertices are assigned the same color. This can be done recursively by observing that each edge belongs to exactly one triangle (otherwise, the endpoints of that edge would have more than one common neighbor, which is not allowed). Moreover, each newly colored vertex always completes a properly colored triangle, ensuring that the coloring remains valid at every step without requiring a fourth color.
  • By the sandwich theorem $\operatorname{\omega}(\mathsf{G})\leq\vartheta(\overline{\mathsf{G}})\leq\operatorname{\chi}(\mathsf{G})$, so $\operatorname{\omega}(\mathsf{G})=\operatorname{\chi}(\mathsf{G})=3$ implies that $\vartheta(\overline{\mathsf{G}})=3$.
  • By Corollary 2.1, where $\mathsf{G}$ is a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,k,1,1)$, we also get
    $\vartheta(\overline{\mathsf{G}})=1+\frac{k}{\sqrt{k-1}}.$
    This leads to a contradiction since, for all $k\geq 3$,
    $(k-2)^{2}>0,$
[…]
    which completes the proof of the theorem.
A Second Alternative Proof of Theorem 3.1: From the point where we get, by contradiction, that $\mathsf{G}$ is a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,k,1,1)$, it is possible to get a contradiction in the following alternative way. This second proof forms a variation of the proof presented in [1, Chapter 44].
Proposition 3.1 (Feasible Parameters of Strongly Regular Graphs).
Let $\mathsf{G}$ be a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$. Then, the following holds:
  1. (1)
    $(n-d-1)\,\mu=d\,(d-\lambda-1)$.
  2. (2)
    $\frac{2d+(n-1)(\lambda-\mu)}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}$ is an integer whose absolute value is at most $n-1$.
  3. (3)
    $6|(nd\lambda)$.
Proof.
This holds by relying on some basic properties of strongly regular graphs [2] as follows.
  • Item 1 is a combinatorial equality for strongly regular graphs. It is obtained by letting the vertices in $\mathsf{G}$ lie in three levels, where an arbitrarily selected vertex is at the root in Level 0, its $d$ neighbors lie in Level 1, and all the others lie in Level 2. The equality in Condition 1 then follows from a double counting argument of the number of edges between Levels 1 and 2.
  • Item 2 holds by the integrality and nonnegativity of the multiplicities $m_{1,2}$ of the second-largest and least eigenvalues of the adjacency matrix, which are respectively given by
    $m_{1,2}=\tfrac{1}{2}\,\Biggl{(}n-1\mp\frac{2d+(n-1)(\lambda-\mu)}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}\,\Biggr{)}.$ (13)
  • Item 3 holds since the number of triangles in $\mathsf{G}$ is given by $\tfrac{1}{6}nd\lambda$. Indeed, every vertex has $d$ neighbors, every pair of adjacent vertices shares exactly $\lambda$ common neighbors (forming a triangle), and each triangle is counted six times due to the six possible permutations of its three vertices.
  • By applying Item 1 in Proposition 3.1 to $\mathsf{G}$ with the parameters $d=k$ and $\lambda=\mu=1$, we get $n=k^{2}-k+1$. This does not lead to a contradiction since summing over all the degrees of the neighbors of an arbitrary vertex $u$ in $\mathsf{G}$ gives $k^{2}$. Then, by the assumption of the theorem that every two vertices have exactly one common neighbor, it follows that the above summation counts every vertex in $\mathsf{G}$ exactly one time, except of $u$ that is counted $k$ times. Hence, indeed $n=k^{2}-k+1$.
  • By Item 2 in Proposition 3.1 with $d=k$ and $\lambda=\mu=1$, we get that $\frac{k}{\sqrt{k-1}}\in\mathbb{N}$. Consequently, $(k-1)|k^{2}\in\mathbb{N}$. Since $k^{2}=(k-1)(k+1)+1$, it follows that $(k-1)|1$, so $k=2$. If $k=2$, the only graph that satisfies the condition of Theorem 3.1 is $\mathsf{G}=\operatorname{\mathsf{K}}_{2}$, which also satisfies the assertion of the theorem. Hence, this argument contradicts the assumption in the proof since it led to the conclusion that $\mathsf{G}$ is a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,k,1,1)$.
4 Spanning Subgraphs of Strongly Regular Graphs
A spanning subgraph is a subgraph obtained by removing some edges from the original graph while preserving all its vertices. Suppose that $\mathsf{G}$ and $\mathsf{H}$ are graphs of the same order, belonging to two given families of strongly regular graphs $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ and $\operatorname{\mathsf{srg}}(n,d^{\prime},\lambda^{\prime},\mu^{\prime})$, respectively (i.e., both graphs have identical orders). The specific structure of these graphs may not be necessarily specified (and recall that there may be various nonisomorphic strongly regular graphs within such a family [3, 2]). Next, we derive a necessary condition for one of these graphs to be a spanning subgraph of the other, relying on the Lovász $\vartheta$-functions of these graphs and their complements.
We next rely on the following simple lemma that relates between the Lovász $\vartheta$-functions of two graphs (not necessarily strongly regular), where one is a spanning subgraph of the other.
Lemma 4.1.
Let $\mathsf{G}$ and $\mathsf{H}$ be undirected and simple graphs with an identical vertex set. If $\mathsf{H}$ is a spanning subgraph of $\mathsf{G}$, then
$\vartheta(\overline{\mathsf{G}})\geq\vartheta(\overline{\mathsf{H}}),$ (14)
$\vartheta(\mathsf{G})\leq\vartheta(\mathsf{H}).$ (15)
Proof.
Inequality (14) follows from (1) since every orthonormal representation of $\overline{\mathsf{G}}$ is also an orthonormal representation of $\overline{\mathsf{H}}$. This holds since if $\{i,j\}\not\in\operatorname{\mathsf{E}}(\overline{\mathsf{H}})$ and $i\neq j$, then $\{i,j\}\in\operatorname{\mathsf{E}}(\mathsf{H})\subseteq\operatorname{\mathsf{E}}(\mathsf{G})$, and therefore $\{i,j\}\not\in\operatorname{\mathsf{E}}(\overline{\mathsf{G}})$; hence, if ${\bf{u}}_{i}^{\mathrm{T}}{\bf{u}}_{j}=0$ for all $\{i,j\}\not\in\operatorname{\mathsf{E}}(\overline{\mathsf{G}})$, then the same also holds for all $\{i,j\}\not\in\operatorname{\mathsf{E}}(\overline{\mathsf{H}})$. Inequality (14) also follows alternatively from the SDP problem in (2) since every feasible solution that corresponds to $\vartheta(\overline{\mathsf{H}})$ is also a feasible solution that corresponds to $\vartheta(\overline{\mathsf{G}})$. Inequality (15) follows from (14) since, by definition, $\mathsf{H}$ is a spanning subgraph of $\mathsf{G}$ if and only if $\overline{\mathsf{G}}$ is a spanning subgraph of $\overline{\mathsf{H}}$. ∎
Remark 4.1.
Let $\mathsf{G}$ and $\mathsf{H}$ be two undirected and simple graphs on $n$ vertices, and suppose that we do not know yet whether $\mathsf{H}$ is a spanning subgraph of $\mathsf{G}$. In light of Corollary 2.2, it follows that if $\mathsf{G}$ and $\mathsf{H}$ are either strongly regular or vertex-transitive graphs, then each of the two inequalities (14) and (15) hold if and only if the other inequality holds. In general, by Corollary 2 of [11], we have $\vartheta(\mathsf{G})\,\vartheta(\overline{\mathsf{G}})\geq n$ and $\vartheta(\mathsf{H})\,\vartheta(\overline{\mathsf{H}})\geq n$, so, if $\mathsf{H}$ is not necessarily a spanning subgraph of $\mathsf{G}$, then the satisfiability of one of the inequalities in (14) and (15) does not necessarily imply the satisfiability of the other inequality.
Before proceeding with the application of Lemma 4.1 to strongly regular graphs, let us consider the problem by only relying on Definition 2.2 of strongly regular graphs. Let $\mathsf{G}$ and $\mathsf{H}$ be strongly regular graphs on $n$ vertices, belonging to the families $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ and $\operatorname{\mathsf{srg}}(n,d^{\prime},\lambda^{\prime},\mu^{\prime})$, respectively. If $\mathsf{H}$ is a spanning subgraph of $\mathsf{G}$, then the following three inequalities hold:
$d>d^{\prime},\quad\lambda\geq\lambda^{\prime},\quad\min\{\lambda,\mu\}\geq\mu^{\prime}.$ (16)
This follows from the fact that deleting edges can only decrease the fixed degree of the vertices in a regular graph. Moreover, since every pair of adjacent vertices in $\mathsf{H}$ are also adjacent in $\mathsf{G}$, the fixed number of common neighbors of any two adjacent vertices in $\mathsf{H}$ must be smaller than or equal to the fixed number of common neighbors of the corresponding vertices in $\mathsf{G}$. Furthermore, the fixed number of common neighbors of any two nonadjacent vertices in $\mathsf{H}$, which are either adjacent or nonadjacent vertices in $\mathsf{G}$, cannot be larger than the minimum between the fixed numbers of common neighbors of adjacent or nonadjacent vertices in $\mathsf{G}$.
We obtain a further necessary condition, which specifies Lemma 4.1 for strongly regular graphs.
Corollary 4.1.
Let $\mathsf{G}$ and $\mathsf{H}$ be strongly regular graphs on $n$ vertices, within the families $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ and $\operatorname{\mathsf{srg}}(n,d^{\prime},\lambda^{\prime},\mu^{\prime})$, respectively. Then, a necessary condition for $\mathsf{H}$ to be a spanning subgraph of $\mathsf{G}$ is that their parameters satisfy the inequality
$\frac{d}{d^{\prime}}\;\frac{\mu^{\prime}-\lambda^{\prime}+\sqrt{(\mu^{\prime}-\lambda^{\prime})^{2}+4(d^{\prime}-\mu^{\prime})}}{\mu-\lambda+\sqrt{(\mu-\lambda)^{2}+4(d-\mu)}}\geq 1.$ (17)
Proof.
The result is obtained by combining Corollary 2.1 (see (9), (10)) and Lemma 4.1 (see (14)), followed by straightforward algebra. ∎
Example 4.1.
Let $\mathsf{G}$ and $\mathsf{H}$ be two strongly regular graphs within the families $\operatorname{\mathsf{srg}}(45,28,15,21)$ and $\operatorname{\mathsf{srg}}(45,22,10,11)$, respectively (by [3], there exist strongly regular graphs with these parameters). The three inequalities in (16) are satisfied, so the question if $\mathsf{H}$ can be a spanning subgraph of $\mathsf{G}$ is left open according to these inequalities. However, by Corollary 4.1, $\mathsf{H}$ cannot be a spanning subgraph of $\mathsf{G}$ since the left-hand side of (17) is equal to $\frac{3\sqrt{5}+1}{11}=0.7007\ldots<1$, so the necessary condition in (17) is violated.
5 Induced Subgraphs of Strongly Regular Graphs
In continuation to Section 4, which is focused on spanning subgraphs of strongly regular graphs, this section derives necessary conditions for a graph to be an induced subgraph of a strongly regular graph belonging to a given family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ with specified parameters (recall that an induced subgraph is obtained by deleting vertices from the original graph, together with all their incident edges). The analysis here relies not only on the Lovász $\vartheta$-function of the strongly regular graph but also on its graph energy.
The graph energy is a spectral graph invariant originally introduced in [7], where its application in chemistry was explored. A comprehensive study of graph energy can be found in the book [10].
Definition 5.1 (Energy of a Graph, [10]).
The energy of a graph $\mathsf{G}$, denoted by $\operatorname{\mathcal{E}}(\mathsf{G})$, is defined to be the sum of the absolute values of the eigenvalues of the adjacency matrix of $\mathsf{G}$, i.e.,
$\operatorname{\mathcal{E}}(\mathsf{G})\triangleq\sum_{i=1}^{n}|\lambda_{i}(\mathsf{G})|.$ (18)
The following lemma relates the energy and the Lovász $\vartheta$-function of an induced subgraph to those of the original graph. It is general and does not assume strong regularity, and we later apply it to strongly regular graphs. It relies on the Cauchy interlacing theorem, which is cited as follows:
Theorem 5.1 (Cauchy interlacing theorem, [4]).
Let $\lambda_{1}\geq\ldots\geq\lambda_{n}$ be the eigenvalues of a symmetric matrix $\mathbf{M}$ and let $\mu_{1}\geq\ldots\geq\mu_{m}$ be the eigenvalues of a principal $m\times m$ submatrix of $\mathbf{M}$ (i.e., a submatrix that is obtained by deleting the same set of rows and columns from $M$). Then, $\lambda_{i}\geq\mu_{i}\geq\lambda_{n-m+i}$ for $i=1,\ldots,m$.
Lemma 5.1.
Let $\mathsf{G}$ and $\mathsf{H}$ be finite, undirected and simple graphs. If $\mathsf{H}$ is an induced subgraph of $\mathsf{G}$, then
$\vartheta(\mathsf{H})\leq\vartheta(\mathsf{G}),$ (19)
$\operatorname{\mathcal{E}}(\mathsf{H})\leq\operatorname{\mathcal{E}}(\mathsf{G}).$ (20)
Proof.
Inequality (19) holds since the vertex set of an induced subgraph $\mathsf{H}$ is a subset of the vertex set of the original graph $\mathsf{G}$, and the adjacency and nonadjacency relations of the unremoved vertices in $\mathsf{H}$ are preserved as compared to those in $\mathsf{G}$. The result then follows from (1) since every orthonormal representation of $\mathsf{G}$ yields an orthonormal representation of $\mathsf{H}$ by taking the subset of the orthonormal vectors $\{{\bf{u}}_{i}\}$ that correspond to the remaining vertices in $\mathsf{H}$.
Inequality (20) follows from the Cauchy interlacing theorem applied to the adjacency matrix of the graph $\mathsf{G}$. Note that the adjacency matrix of the induced subgraph $\mathsf{H}$ is obtained from the adjacency matrix of $\mathsf{G}$ by deleting the rows and columns that correspond to the deleted vertices in the subgraph $\mathsf{H}$. ∎
Remark 5.1.
[…]
We next rely on the spectral properties of strongly regular graphs, which are listed as follows, to derive a closed-form expression of their energy. Note that if $\mathsf{G}$ is a strongly regular graph that belongs to the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$, then $\mathsf{G}$ is connected if and only if $\mu>0$. In that case, the diameter of $\mathsf{G}$ is equal to 2 (a complete graph is excluded from the family of strongly regular graphs).
Theorem 5.2 (The adjacency spactra of strongly regular graphs, [2]).
The following properties hold:
  1. (1)
    A strongly regular graph has three distinct eigenvalues.
  2. (2)
    Let $\mathsf{G}$ be a connected strongly regular graph that belongs to the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$. Then, the largest eigenvalue of its adjacency matrix is $\lambda_{1}(\mathsf{G})=d$ with multiplicity 1, and the other two distinct eigenvalues of its adjacency matrix are given by
    $p_{1,2}=\tfrac{1}{2}\,\Biggl{(}\lambda-\mu\pm\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}\,\Biggr{)},$ (21)
    with the respective multiplicities in (13).
  3. (3)
    A connected regular graph with exactly three distinct eigenvalues is strongly regular.
  4. (4)
    Strongly regular graphs for which $2d+(n-1)(\lambda-\mu)\neq 0$ have integral eigenvalues and their respective multiplicities are distinct.
  5. (5)
    A connected regular graph is strongly regular if and only if it has three distinct eigenvalues, where the largest eigenvalue is of multiplicity 1.
By Theorem 5.2, the spectrum of a strongly regular graph has exactly three distinct eigenvalues. A closed-form expression for the energy of strongly regular graphs with general parameters does not seem, however, to appear in the literature (see, e.g., [2, 8, 9, 10]), and it is next introduced for the continuation of the analysis in this section.
Lemma 5.2 (The energy of strongly regular graphs).
Let $\mathsf{G}$ be a strongly regular graph that belongs to the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ with $\mu>0$. Then, its energy is given by
$\operatorname{\mathcal{E}}(\mathsf{G})=d\,\Biggl{(}1+\frac{2(n-d)+\lambda-\mu}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}\Biggr{)}.$ (22)
Proof.
By the spectral properties of strongly regular graphs, the adjacency matrix of a strongly regular graph has three distinct eigenvalues. The largest eigenvalue of $\mathsf{G}$ is equal to $d$, the second largest eigenvalue of $\mathsf{G}$ is $\lambda_{2}(\mathsf{G})=p_{1}$ with multiplicity $m_{1}$, and the smallest eigenvalue of $\mathsf{G}$ is $\lambda_{\min}(\mathsf{G})=p_{2}$ with multiplicity $m_{2}$ (see (21) and (13)). Consequently, the energy of $\mathsf{G}$ is given by
$\operatorname{\mathcal{E}}(\mathsf{G})$ $=d+m_{1}|p_{1}|+m_{2}|p_{2}|$
$=d+\tfrac{1}{2}\,\Biggl{(}\lambda-\mu+\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}\,\Biggr{)}\cdot\tfrac{1}{2}\,\Biggl{(}n-1-\frac{2d+(n-1)(\lambda-\mu)}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}\,\Biggr{)}$
$\hskip 5.69046pt+\tfrac{1}{2}\,\Biggl{(}\mu-\lambda+\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}\Biggr{)}\cdot\tfrac{1}{2}\,\Biggl{(}n-1+\frac{2d+(n-1)(\lambda-\mu)}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}\,\Biggr{)}$
$=d\,\Biggl{(}1+\frac{2(n-d)+\lambda-\mu}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}\Biggr{)},$ (23)
where the last equality in (23) can be verified by straightforward algebra, and relying on the connection between the four parameters of a strongly regular graph as given in Item 1 of Proposition 3.1. ∎
Example 5.1.
The adjacency spectrum of the Petersen graph is the multiset $\{3^{1},1^{5},(-2)^{4}\}$, where superscripts indicate the multiplicities of the eigenvalues. Hence, its graph energy is 16. Since it is a strongly regular graph with parameters $\operatorname{\mathsf{srg}}(10,3,0,1)$, their substitution into (22) confirms this result.
In the context of Lemma 5.2, it is worth noting the following result on maximal energy graphs, which combines Theorems 5.9 and 5.10 of [10].
Theorem 5.3 (Maximal energy graphs, [8, 9]).
Let $\mathsf{G}$ be a graph on $n$ vertices. Then,
$\operatorname{\mathcal{E}}(\mathsf{G})\leq\tfrac{1}{2}n(\sqrt{n}+1),$ (24)
with equality if $\mathsf{G}$ is a strongly regular graph with parameters
$\operatorname{\mathsf{srg}}\Bigl{(}n,\tfrac{1}{2}(n+\sqrt{n}),\tfrac{1}{4}(n+2\sqrt{n}),\tfrac{1}{4}(n+2\sqrt{n})\Bigr{)}.$ (25)
Furthermore, strongly regular graphs with parameters as in (25) exist if
  1. (1)
    $n=4^{p}q^{4}$ with $p,q\in\mathbb{N}$;
  2. (2)
    $n=4^{p+1}q^{2}$ with $p\in\mathbb{N}$ and $14q-$ 1 is a prime power, or $12q-$ 1 is a prime power, or $q$ is a square of an integer, or $q<167$.
Corollary 5.1.
Let $\mathsf{H}$ be an induced subgraph of a strongly regular graph $\mathsf{G}$ that belongs of the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ with $\mu>0$. Then, the following inequalities hold:
  1. (1)
    $\vartheta(\mathsf{H})\leq\dfrac{n\,(t+\mu-\lambda)}{2d+t+\mu-\lambda},$ (26)
    $\operatorname{\mathcal{E}}(\mathsf{H})\leq d\,\biggl{(}1+\frac{2(n-d)+\lambda-\mu}{t}\biggr{)},$ (27)
    where $t\triangleq\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}$, as given in (10).
  2. (2)
    Specifically, if $\mathsf{H}$ belongs to the family $\operatorname{\mathsf{srg}}(n^{\prime},d^{\prime},\lambda^{\prime},\mu^{\prime})$, where $n^{\prime}\leq n$, $d^{\prime}\leq d$, $\lambda^{\prime}\leq\lambda$, and $\nu^{\prime}\leq\nu$, then
    $\dfrac{n^{\prime}}{n}\cdot\dfrac{t^{\prime}+\mu^{\prime}-\lambda^{\prime}}{t+\mu-\lambda}\cdot\dfrac{2d+t+\mu-\lambda}{2d^{\prime}+t^{\prime}+\mu^{\prime}-\lambda^{\prime}}\leq 1,$ (28)
    $\frac{td^{\prime}}{t^{\prime}d}\cdot\frac{t^{\prime}+2(n^{\prime}-d^{\prime})+\lambda^{\prime}-\mu^{\prime}}{t+2(n-d)+\lambda-\mu}\leq 1,$ (29)
    where $t^{\prime}\triangleq\sqrt{(\lambda^{\prime}-\mu^{\prime})^{2}+4(d^{\prime}-\mu^{\prime})}$.
Proof.
Inequalities (26)–(29) follow from Lemma 5.1, combined with its specialization according to Corollary 2.1 and Lemma 5.2. ∎
Example 5.2 (Induced subgraphs of triangular graphs).
The family of triangular graphs consists of regular graphs obtained as the line graphs of complete graphs on at least three vertices. Let $T_{\ell}$ denote the line graph of the complete graph $\operatorname{\mathsf{C}}_{\ell}$, for all $\ell\geq 3$. This results in a strongly regular graph belonging to the family $\operatorname{\mathsf{srg}}(\tfrac{1}{2}\ell(\ell-1),2\ell-4,\ell-2,4)$ for all $\ell\geq 4$ (see [2, Section 1.1.7]). For $\ell\geq 4$ with $\ell\neq 8$, these connected, strongly regular graphs are unique for their respective parameters. Their uniqueness further implies that they are determined by their spectra (following from Theorem 34 of [16]).
Now, let $\mathsf{H}$ be a strongly regular graph that belongs to the family $\operatorname{\mathsf{srg}}(n^{\prime},d^{\prime},\lambda^{\prime},\mu^{\prime})$ for some specified (feasible) parameters. We consider whether $\mathsf{H}$ can appear as an induced subgraph of the triangular graph $T_{\ell}$, as a function of $\ell\geq 4$. Applying Item 2 of Corollary 5.1, if either inequality (28) or (29) is violated, then no strongly regular graph in the family $\operatorname{\mathsf{srg}}(n^{\prime},d^{\prime},\lambda^{\prime},\mu^{\prime})$ (including $\mathsf{H}$) can be an induced subgraph of $T_{\ell}$. This leads to a general conclusion for those values of $\ell$ where either inequality fails. Applying Item 2 of Corollary 5.1 to $T_{\ell}$ with the parameters:
$n=\tfrac{1}{2}\ell(\ell-1),\quad d=2(\ell-2),\quad\lambda=\ell-2,\quad\mu=4,$ (30)
yields, by (10),
$t=\ell-2,\quad\forall\,\ell\geq 4.$ (31)
With straightforward algebra, the inequalities (28) and (29) simplify to
[…]
where $t^{\prime}\triangleq\sqrt{(\lambda^{\prime}-\mu^{\prime})^{2}+4(d^{\prime}-\mu^{\prime})}$. If, for given parameters $(n^{\prime},d^{\prime},\lambda^{\prime},\mu^{\prime})$, at least one of these inequalities is violated for a specific $\ell\geq 4$, then $\mathsf{H}$ cannot be an induced subgraph of $T_{\ell}$.
Example 5.3 (Continuation to Example 5.2: the Gewirtz graph).
Consider the Gewirtz graph, denoted by $\mathsf{H}$, which is a connected, triangle-free, strongly regular graph belonging to the family $\operatorname{\mathsf{srg}}(56,10,0,2)$ (see [2, Section 10.20]). To clarify, since $\mu=2>0$, the graph $\mathsf{H}$ is connected, and $\lambda=0$ confirms that it is triangle-free. Uniqueness in this parameter family ensures it is determined by its spectrum [16]. Specializing (32) for $\mathsf{H}$, the first inequality is violated if and only if $\ell\leq 31$ (it dominates the second, energy-based, inequality in (32)). Hence, $\mathsf{H}$ cannot be an induced subgraph of $T_{\ell}$ for all $\ell\leq 31$. For $\ell\leq 11$, this result is trivial since $\tfrac{1}{2}\ell(\ell-1)<56$, meaning that the order of $T_{\ell}$ is too small to contain $\mathsf{H}$ as a subgraph. However, for $\ell\geq 32$, this question remains open, as both inequalities in (32) hold.